# Better bootstrap $t$ confidence intervals for the mean

### v1

@@ 1 Introduction
This article is about nonparametric approximate confidence intervals (ACIs) for the mean of a real random variable. There are two leading bootstrap solutions for this problem. One is the bootstrap $t$ method and the other is the bias corrected accelerated (BCa) method. These are both well understood theoretically in the limit as the sample size $n$ of an IID sample diverges to infinity. Under mild conditions on the sampling distribution, both are second order accurate with both one and two sided coverage errors decaying as $O(1/n)$. Alternative methods that are not second order accurate typically have one sided coverage errors of $O(1/\sqrt{n})$ with two sided errors of $O(1/n)$ (Hall,, 1988).
@@
We can eliminate the zero weights by merging the Bayesian bootstrap (Rubin,, 1981) with the bootstrap $t$. That provides asymptotically valid confidence intervals, though it lacks second order accuracy for one-sided confidence intervals. Unfortunately, the Bayesian bootstrap version of the bootstrap $t$ intervals does not have very good small sample coverage in the cases examined here. Perhaps the good coverage properties of the original bootstrap $t$ are strongly dependent on those small weights. A next step is to then to replace the unit exponential weights of the Bayesian bootstrap by weights from a distribution that has a singular probability density at the origin in order to get near zero weights that remain positive. Strictly positive weights keep the resampled variance positive unless all $n$ observations are equal. A positive resample variance then keeps the resampled $t$ value finite. One such weight distribution is Beta($1/2,3/2$). We show that it has a second order accuracy property as $n\to\infty$, so its small sample advantages do not conflict with asymptotic performance.
@@ 2 Background and notation
We suppose throughout that $x_{1},\dots,x_{n}$ are sampled IID from a distribution $F$ with mean $\mu$ and variance $\sigma^{2}\in(0,\infty)$. A confidence interval for $\mu$, at level $\alpha\in(0,1)$, is a pair of random variables $L$ and $U$, given as functions of $x_{1},\dots,x_{n}$ for which $\Pr(L\leqslant\mu\leqslant U)=\alpha$. A customary default choice is $\alpha=0.95$. For one-sided intervals, we select $0<\alpha_{1}<\alpha_{2}<1$ and seek random $L$ and $U$ with $\Pr(\mu<L)=\alpha_{1}$ and $\Pr(\mu\leqslant U)=\alpha_{2}$ so that $\Pr(L\leqslant\mu\leqslant U)=\alpha_{2}-\alpha_{1}$. For equal one-sided coverage errors we take $\alpha_{1}=(1-\alpha)/2$ and $\alpha_{2}=1-\alpha_{1}=(1+\alpha)/2$. Then the customary values are $\alpha_{1}=0.025$ and $\alpha_{2}=0.975$. While $0.95$ is not a very high confidence level, it is already a challenge to get close to this coverage in small samples, even for two-sided coverage.
@@
When $F$ does not belong to a known parametric family and $x$ does not satisfy some special side condition (such as known bounds or symmetry) then we must rely instead on approximate confidence intervals (ACIs). These are usually based on the central limit theorem where $\sqrt{n}(\bar{x}-\mu)/\sigma\stackrel{{\scriptstyle\cdot}}{{\sim}}\mathcal{N}(0,1)$ for ‘large enough’ $n$. There is no sharp distinction where we can say that $n$ is large enough while n1n-1 isn’t. A very common rule of thumb is to consider $n\geqslant 30$ large enough. Kufs, (2010) looks into the history of this rule of thumb. Miller, (1986) even suggests that $n\geqslant 10$ may be enough, if one can accept mild skewness in $\bar{x}$.
@@ 2.2 Bootstrap $t$
Hall and Mammen, (1994) consider several ways to define a bootstrap $t$. When their weights $N_{ni}^{*}$ are normalized to sum to $n$, like $nw_{i}$ here, then their $\hat{\sigma}^{*2}$ on page 2017 reduces to $\hat{\sigma}^{*2}$ above. They refer $(\bar{x}-\mu)/\hat{\sigma}$ to $(\bar{x}^{*}-\bar{x}/\hat{\sigma}^{*}$ which is the same bootstrap approximation we use here. Mason and Newton, (1992) study a bootstrap $t$ method where $\hat{\sigma}^{*}$ is multiplied by a standard deviation of the resampling weights. The multinomial bootstrap $t$ is not then a special case of their quantity. Csörgő et al., (2014) consider this same statistic as well as $\sqrt{\sum_{i=1}^{n}v_{i}}(\bar{x}^{*}-\bar{x})/\hat{\sigma}^{*}$. That latter statistic does generalize $t^{*}$. They show consistency but do not consider convergence rates.
@@ 2.4 Bootstrap as reweighting
There has not been much attention paid to the kurtosis of the pseudo-counts. One exception is Guillou, (1999) who gives a kurtosis condition that makes the one-sided coverage errors $O(n^{-3/2})$ for studentized statistics like we study here. In other words, he can obtain third order accuracy. In addition to a Cramer condition, Guillou requires the sample skewness of $v_{i}$ to be $1+O_{p}(n^{-1/2})$ and also requires the sample kurtosis of the $v_{i}$ to satisfy a relation with the sample kurtosis of the $x_{i}$.
@@ 4.2 Elicitation and the interval score
$\theta_{2}-\theta_{1}+\frac{1_{x<\theta_{1}}}{\alpha_{1}}(\theta_{1}-x)+\frac{1_{x>\theta_{2}}}{\alpha_{1}}(x-\theta_{2}).$ (13)
The above derivation uses $1_{x\leqslant\theta_{1}}(\theta_{1}-x)=1_{x<\theta_{1}}(\theta_{1}-x)$. Equation (13) is equation (43) of Gneiting and Raftery, (2007). Their $\alpha$ is our $1-\alpha$, so for the default coverage setting, our $1/\alpha_{1}=40$ is their $2/\alpha=40$.
When scoring confidence intervals $[L,U]$ for the mean $\mu$, if we reverse the notion of what is random, as if $\mu$ were random with estimated quantiles $L$ and $U$, then the interval score (13) yields
@@
for $x_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}F$. By the result of Bahadur and Savage, (1956) we know that no non-trivial such $T_{n}$ exists. Perhaps we should not expect to elicit a statistical quantity that does not exist. Some quantities that are closely related to confidence intervals are known to not be elicitable. Frongillo and Kash, (2014) call the interval $(a,b)$ a 90% confidence interval for $x$ if $\int_{a}^{b}p(x)\,\mathrm{d}x=0.9$. That does not correspond to our usage of the term ‘confidence interval’. It is a prediction interval. They go on to show that the entire set of such pairs $(a,b)$ is not elicitable, nor is the pair with the smallest width $\omega=b-a$, nor is $\omega$ itself elicitable. From here on, we assume that there is no consistent score function for non-trivial confidence intervals.
@@ 4.4 The $F_{1}$ score
We can draw an analogy between binary classification and confidence intervals. The true value $\mu$ is the one real number that we wish to recover. That makes coverage probability the analogue of recall. Taking $\tilde{r}=\max(\Pr(L\leqslant\mu\leqslant U),0.95)$ allows us to incorporate a specific nominal coverage level, which we were not able to do with the interval score. This criterion is designed to reward absolute coverage (up to a limit) instead of penalizing coverage error.
@@ 4.6 Illustration
Method Cover Length Length2 GR95 GR99 $\tilde{F}_{1}$
@@
Table 3: This table describes 9595% ACIs for a sample of $n=10$ draws from $\mathrm{Exp}(1)$. All the methods are weighted bootstrap $t$ except for BCa and Student’s $t$. The columns are mean coverage, mean interval length, mean squared interval length, two Gneiting-Raftery interval scores, one at level $0.95$ and one at level $0.99$, and $\tilde{F}_{1}$ from (21).
If we use the Gneiting-Raftery interval score for level $0.95$ then the best score comes from a bootstrap $t$ using Rubin’s Bayesian bootstrap, i.e., exponential weights. It does not seem reasonable that a method with 89.3589.35% coverage should be best when other methods attain 93% to 96% coverage. Those other methods use longer intervals, but they do not appear to have drastically greater length.
Sometimes experts making a prediction might make a prediction that is more cautious than their true belief if the score function is not proper. In a setting with no proper score function like ACIs appear to be, we might judge that the 95% interval score is too lenient on coverage (versus length) and compensate by scoring nominally 95% ACIs by a higher coverage goal such as 99%. By that measure we would favor the bootstrap $t$ with power weights. It raises the coverage from just over 89% to 94% while increasing the mean length by a factor of $1.62/1.23\doteq 1.32$ and the mean squared length by $3.32/1.84\doteq 1.80$.
Poisson weights attain essentially the same score as power weights do here. They do so by getting credit for coverage that goes beyond even 95%. While it may or may not be reasonable to penalize a method for covering more than the nominal level, it does not seem reasonable to credit it for covering more than the nominal level. This breaks the interval score tie in favor of the power bootstrap $t$.
@@ 5.1 Asymptotic coverage
The coverage errors for 95% confidence intervals are $\Pr(L\leqslant\mu\leqslant U)-0.95$, so negative values indicate undercoverage. The asymptotic formulas involve the skewness $\gamma=\mathbb{E}((x-\mu)^{3})/\sigma^{3}$ and the kurtosis $\kappa=\mathbb{E}((x-\mu)^{4})/\sigma^{4}-3$. The coverage errors for two sided 95% confidence intervals are
@@ 5.2 Computations
Each sampling distribution was represented in R=10, [^0] R=10{,}000 data sets of sizes $2\leqslant n\leqslant 20$. All of the ACI methods were given the same 10,00010{,}000 samples of each size so they all received the same problem difficulty. Furthermore the samples of size $n+1$ always contained the points from the samples of size $n$. That technique gives more precise estimates of how length or coverage vary with $n$. Every bootstrap was based on $B=2000$ resamplings.
@@ 5.4 Skewed/exponential data
Skewness is one of the main reasons to switch from Student’s $t$ distribution to some kind of bootstrap distribution. Our archetypal example for mildly skewed data has $x_{i}\sim\mathrm{Exp}(1)$. It has skewness $2$. There is no universal rule for what amount of skewness is mild and what is severe. Some authors consider skewness in [1,1][-1,1] to be mild and skewness $2$ to already be severe. Others consider skewness of $2$ to not be extreme.
@@ Acknowledgments
This work was supported by the National Science Foundation under grant DMS-2152780. I thank Tilmman Gneiting, Enno Mammen and Lisa Hofer for helpful discussions.

### v2

@@ 1 Introduction
This article is about nonparametric approximate confidence intervals (ACIs) for the mean of a real random variable. There are two leading bootstrap solutions for this problem. One is the bootstrap $t$ method and the other is the bias corrected accelerated (BCa) method. These are both well understood theoretically in the limit as the sample size $n$ of an IID sample diverges to infinity. Under mild conditions on the sampling distribution, both are second order accurate with both one and two-sided coverage errors decaying as $O(1/n)$. Alternative methods that are not second order accurate typically have one-sided coverage errors of $O(1/\sqrt{n})$ with two-sided errors of $O(1/n)$ (Hall,, 1988).
@@
We can eliminate the zero weights by merging the Bayesian bootstrap (Rubin,, 1981) with the bootstrap $t$. That provides asymptotically valid confidence intervals, though it lacks second order accuracy for one-sided confidence intervals. Unfortunately, the Bayesian bootstrap version of the bootstrap $t$ intervals does not have very good small sample coverage in the cases examined here. Perhaps the good coverage properties of the original bootstrap $t$ are strongly dependent on those small weights. A next step is to then to replace the unit exponential weights of the Bayesian bootstrap by weights from a distribution that has a singular probability density at the origin in order to get near zero weights that remain positive. Strictly positive weights keep the resampled variance positive unless all $n$ observations are equal. That positive resample variance then keeps the resampled $t$ value finite. One such weight distribution is Beta($1/2,3/2$). We show that it has a second order accuracy property as $n\to\infty$, so its small sample advantages do not conflict with asymptotic performance.
@@ 2 Background and notation
We suppose throughout that $x_{1},\dots,x_{n}$ are sampled IID from a distribution $F$ with mean $\mu$ and variance $\sigma^{2}\in(0,\infty)$. A two-sided confidence interval for $\mu$, at level $\alpha\in(0,1)$, is a pair of random variables $L$ and $U$, given as functions of $x_{1},\dots,x_{n}$ for which $\Pr(L\leqslant\mu\leqslant U)=\alpha$. We always have $L\leqslant U$. A customary default choice is $\alpha=0.95$. While $0.95$ is not a very high confidence level, it is already a challenge to get close to this coverage in small samples, even for two-sided coverage.
The two-sided intervals we consider are formed from a pair of one-sided intervals. We select $0<\alpha_{1}<\alpha_{2}<1$ and seek random $L$ and $U$ with $\Pr(\mu<L)=\alpha_{1}$ and $\Pr(\mu\leqslant U)=\alpha_{2}$ so that $\Pr(L\leqslant\mu\leqslant U)=\alpha_{2}-\alpha_{1}$. For equal one-sided coverage errors we take $\alpha_{1}=(1-\alpha)/2$ for $\alpha>1/2$ and $\alpha_{2}=1-\alpha_{1}=(1+\alpha)/2$. Then the customary values are $\alpha_{1}=0.025$ and $\alpha_{2}=0.975$.
@@
When $F$ does not belong to a known parametric family and $x$ does not satisfy some special side condition (such as known bounds or symmetry) then we must rely instead on approximate confidence intervals (ACIs). These are usually based on the central limit theorem where $\sqrt{n}(\bar{x}-\mu)/\sigma\stackrel{{\scriptstyle\cdot}}{{\sim}}\mathcal{N}(0,1)$ for ‘large enough’ $n$. There is no sharp distinction where we can say that $n$ is large enough while n1n-1 isn’t. A very common rule of thumb is to consider $n\geqslant 30$ large enough. Kufs, (2010) looks into the history of this rule of thumb. Miller, (1986) even suggests that $n\geqslant 10$ may be enough, if one can accept mild skewness in $\bar{x}$. On the other hand Hesterberg, (2015) has stringent conditions (one-sided coverages within $0.0025$ of $0.025$) where Student’s $t$ needs $n\geqslant 4815$ for $x_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Exp}(1)$. Based on some simulations, Boos and Hughes-Oliver, (2000) suggest $n\geqslant(5.66\gamma)^{2}$ to get nominal 95% intervals Student’s $t$ intervals with two-sided coverage at least 9494%, when $x_{i}$ have skewness $\gamma$. Kohavi et al., (2014) find that for some highly skewed electronic commerce data this guideline requires $n\geqslant 114{,}000$. These higher numbers provide good motivation for further investigation of confidence intervals.
@@ 2.2 Bootstrap $t$
Hall and Mammen, (1994) consider several ways to define a bootstrap $t$. When their weights $N_{ni}^{*}$ are normalized to sum to $n$, like $nw_{i}$ here, then their $\hat{\sigma}^{*2}$ on page 2017 reduces to $\hat{\sigma}^{*2}$ above. They refer $(\bar{x}-\mu)/\hat{\sigma}$ to $(\bar{x}^{*}-\bar{x})/\hat{\sigma}^{*}$ which is the same bootstrap approximation we use here. Mason and Newton, (1992) study a bootstrap $t$ method where $\hat{\sigma}^{*}$ is multiplied by a standard deviation of the resampling weights. The multinomial bootstrap $t$ is not then a special case of their quantity. Csörgő et al., (2014) consider this same statistic as well as $\sqrt{\sum_{i=1}^{n}v_{i}}(\bar{x}^{*}-\bar{x})/\hat{\sigma}^{*}$. That latter statistic does generalize $t^{*}$. They show consistency but do not consider convergence rates.
@@ 2.4 Bootstrap as reweighting
There has not been much attention paid to the kurtosis of the pseudo-counts. One exception is Guillou, (1999) who gives a kurtosis condition that makes the one-sided coverage errors $O(n^{-3/2})$ for studentized statistics like we study here. In other words, she can obtain third order accuracy. In addition to a Cramer condition, Guillou requires the sample skewness of $v_{i}$ to be $1+O_{p}(n^{-1/2})$ and also requires the sample kurtosis of the $v_{i}$ to satisfy a relation with the sample kurtosis of the $x_{i}$.
@@ 4.2 Elicitation and the interval score
$\theta_{2}-\theta_{1}+\frac{1_{x<\theta_{1}}}{\alpha_{1}}(\theta_{1}-x)+\frac{1_{x>\theta_{2}}}{\alpha_{1}}(x-\theta_{2})$ (13)
due to Winkler, (1972). The above derivation uses $1_{x\leqslant\theta_{1}}(\theta_{1}-x)=1_{x<\theta_{1}}(\theta_{1}-x)$. Equation (13) is equation (43) of Gneiting and Raftery, (2007). Their $\alpha$ is our $1-\alpha$, so for the default coverage setting, our $1/\alpha_{1}=40$ is their $2/\alpha=40$.
When scoring confidence intervals $[L,U]$ for the mean $\mu$, if we reverse the notion of what is random, as if $\mu$ were random with estimated quantiles $L$ and $U$, then the Winkler score (13) yields
@@
for $x_{i}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}F$. By the result of Bahadur and Savage, (1956) we know that no non-trivial such $T_{n}$ exists. Perhaps we should not expect to elicit a statistical quantity that does not exist. Some quantities that are closely related to confidence intervals are known to not be elicitable. Frongillo and Kash, (2014) call the interval $(a,b)$ a 90% confidence interval for $x$ if $\int_{a}^{b}p(x)\,\mathrm{d}x=0.9$. That does not correspond to our usage of the term ‘confidence interval’. It is a prediction interval. They go on to show that the entire set of such pairs $(a,b)$ is not elicitable, nor is the pair with the smallest width $\omega=b-a$, nor is $\omega$ itself elicitable. For more in this direction, see Brehmer and Gneiting, (2021). From here on, we assume that there is no consistent score function for non-trivial confidence intervals.
@@ 4.4 The $F_{1}$ score
We can draw an analogy between binary classification and confidence intervals. The true value $\mu$ is the one real number that we wish to recover. That makes coverage probability the analogue of recall. Taking $\tilde{r}=\max(\Pr(L\leqslant\mu\leqslant U),0.95)$ allows us to incorporate a specific nominal coverage level, which we were not able to do with the Winkler score. This criterion is designed to reward absolute coverage (up to a limit) instead of penalizing coverage error.
@@ 4.6 Illustration
Method Cover Length Length2 Win95 Win99 $\tilde{F}_{1}$
@@
Table 3: This table describes 9595% ACIs for a sample of $n=10$ draws from $\mathrm{Exp}(1)$. All the methods are weighted bootstrap $t$ except for BCa and Student’s $t$. The columns are mean coverage, mean interval length, mean squared interval length, two Winkler scores, one at level $0.95$ and one at level $0.99$, and $\tilde{F}_{1}$ from (21).
If we use the Winkler score for level $0.95$ then the best score comes from a bootstrap $t$ using Rubin’s Bayesian bootstrap, i.e., exponential weights. It does not seem reasonable that a method with 89.3589.35% coverage should be best when other methods attain 93% to 96% coverage. Those other methods use longer intervals, but they do not appear to have drastically greater length.
Sometimes experts making a prediction might make a prediction that is more cautious than their true belief if the score function is not proper. In a setting with no proper score function like ACIs appear to be, we might judge that the 95% Winkler score is too lenient on coverage (versus length) and compensate by scoring nominally 95% ACIs by a higher coverage goal such as 99%. By that measure we would favor the bootstrap $t$ with power weights. It raises the coverage from just over 89% to 94% while increasing the mean length by a factor of $1.62/1.23\doteq 1.32$ and the mean squared length by $3.32/1.84\doteq 1.80$.
Poisson weights attain essentially the same score as power weights do here. They do so by getting credit for coverage that goes beyond even 95%. While it may or may not be reasonable to penalize a method for covering more than the nominal level, it does not seem reasonable to credit it for covering more than the nominal level. This breaks the Winkler score tie in favor of the power bootstrap $t$.
@@ 5.1 Asymptotic coverage
The coverage errors for 95% confidence intervals are $\Pr(L\leqslant\mu\leqslant U)-0.95$, so negative values indicate undercoverage. The asymptotic formulas involve the skewness $\gamma=\mathbb{E}((x-\mu)^{3})/\sigma^{3}$ and the kurtosis $\kappa=\mathbb{E}((x-\mu)^{4})/\sigma^{4}-3$. The coverage errors for two-sided 95% confidence intervals are
@@ 5.2 Computations
Each sampling distribution was represented in R=10, [^0] R=10{,}000 data sets of sizes $2\leqslant n\leqslant 20$. All of the ACI methods were given the same 10,00010{,}000 samples of each size so they all received the same problem difficulty. Furthermore the samples of size $n+1$ always contained the points from the samples of size $n$. That technique gives more precise estimates of how length or coverage vary with $n$. For R=10, [^0] R=10{,}000 replicates, and a method with coverage probability $0.95$, the observed number of of coverages has $0.005$ quantile 9443 and $0.995$ quantile 9555. Then the coverage proportion has about 99.999.9% probability of being within $0.57$ percentage points of nominal. For a method with coverage 8080% comparable intervals are roughly $0.80\pm 0.01$. Every bootstrap was based on $B=2000$ resamplings.
@@ 5.4 Skewed/exponential data
Skewness is one of the main reasons to switch from Student’s $t$ distribution to some kind of bootstrap distribution. Our archetypal example for mildly skewed data has $x_{i}\sim\mathrm{Exp}(1)$. It has skewness $2$. There is no universal rule for what amount of skewness is mild and what is severe. Some authors consider skewness in [1,1][-1,1] to be mild and skewness $2$ to already be severe. Others consider skewness of $2$ to not be extreme. Where Hesterberg, (2015) recommends $n\geqslant 4815$ for Student’s $t$ test he finds that the bootstrap $t$ with $n\geqslant 101$ meets his stringent conditions.
@@ Acknowledgments
This work was supported by the National Science Foundation under grant DMS-2152780. I thank Tilmman Gneiting, Enno Mammen, Lisa Hofer and Tim Hesterberg for helpful discussions.